Write the reactions for the dissolution of the following compounds in aqueous medium: Mg(OH)2 (s) AgI...

  1. Write the reactions for the dissolution of the following compounds in aqueous medium:
  1. Mg(OH)2 (s)
  2. AgI (s)
  3. Ba3(PO4)2 (s)
Answer #1

a, Mg(OH)2 (8) = Mght (ag) + 20H+ (aq) b. Ag I(3) = Agt (aq) + I + (aq) e. Bag(PO4)2 (s) = 3B&+(aq) + 2 POMP Cag)